OpenStudy (lgbasallote):

Openstudy user L's medal count is 112 less than the medal count of OpenStudy user P. OpenStudy user P's accuracy (medals received-over-questions answered) is 79.85%. OpenStudy user L's accuracy is 62.41% who has answered more questions? can someone tell me how to express this in algebra??

Parth (parthkohli):

lol P is me

OpenStudy (lgbasallote):

hey

OpenStudy (lgbasallote):

lol you sent faster than me

Parth (parthkohli):

You have answered more xP

OpenStudy (lgbasallote):

and no it aint you..feeler :p

OpenStudy (lgbasallote):

so how to express this algebraically @ParthKohli ?

Parth (parthkohli):

Your accuracy is less. I think it has to do something with inverse relation.

Parth (parthkohli):

The lesser the accuracy, the more questions answered.

OpenStudy (lgbasallote):

lol..is that a legit logic o.O

Parth (parthkohli):

Yeah

OpenStudy (lgbasallote):

well my book requires a more legit solution :p

Parth (parthkohli):

I compared the medal count and I have exactly 112 more :P

OpenStudy (lgbasallote):

hah lol

OpenStudy (lgbasallote):

proves nothing

OpenStudy (kropot72):

Let P's medal count = x L's medal count = x - 112 P's number of questions = x/0.7985 L's number of questions = (x - 112)/0.6241 = (x/0.6241) -179

OpenStudy (lgbasallote):

so i guess L has more right?

OpenStudy (kropot72):

Dunnow :(